A +13 nC charge is located at the origin.

A-

What is the electric field at the position (x2,y2)=(−5.0cm,5.0cm)? Write electric field vector in component form.

B-

What is the electric field at the position

(x3,y3)=(−5.0cm,−5.0cm)? Write electric field vector in component form.
Answer #1

Given data

q = 13 nc = 13*10^-9 C

A) electric field at the given point,

E = k*q/d^2

= 9*10^9*13*10^-9/(0.05^2 + 0.05^2)

= 23400 N/c

Ex = -E*cos(45) = -23400*cos(45) = -16546.3 N/c

Ey = E*sin(45) = 23400*sin(45) = 16546.3 N/c

so at (-5cm,5cm)

E = (-16546.3 i + 16546.3 j ) N/c

B)

at (-5cm,-5cm)

E = (-16546.3 i - 16546.3 j ) N/c